OpenStudy (anonymous):

An astronaut who weighed 170 pounds on earth found that his weight on jupiter was 400 pounds. he captured an alien that weighed 120 pounds on jupiter. what is the alien's weight on earth?

OpenStudy (anonymous):

You can do proportions here\[\frac{ earth }{ jupiter }=>\frac{ 170 }{ 400 }=\frac{ x }{ 120 }\]So cross multiply to find x.